OpenStudy (anonymous):

Find derivative of e^2x/3

OpenStudy (johnweldon1993):

Remember the chain rule \[\large \frac{d}{dx} e^{ax} = \frac{d}{dx} e^x \times \frac{d}{dx} ax\]

OpenStudy (anonymous):

So ax =2x/3?